If f(x) =2x-5 and g(x)=2, find
a) f(1)
b) g(2)
c) gf(1)
d) fg(2)

Answer by ikleyn(52778)

.
If f(x) =2x-5 and g(x)=2, find
a) f(1)
b) g(2)
c) gf(1)
d) fg(2)
------------------------------

a) Please do it yourself. Simply substitute x = 1 into f(x).


b) Please do it yourself. Simply substitute x = 2 into g(x).


c) The symbol gf indicates here the composition function (gof)(x).

   It is  gf(x) = (gof)(x) = g(f(x)) == 2  due to the definition of g.

   So,  gf(1) = 2. 


d) The symbol fg indicates here the composition function (fog)(x).

   It is  fg(x) = (fog)(x) = f(g(x)) = f(2) = 2*2 - 5 = 4 - 1 = -1, the constant value (does not depend on x).

   So,  fg(1) = -1.